OpenStudy (kaywee):

A bag contains 5 blue marbles, 2 black marbles, and 3 red marbles. A marble is randomly drawn from the bag. The probability of not drawing a black marble is . The probability of drawing a red marble is . NextReset

OpenStudy (zzr0ck3r):

The probability of not drawing a black marble from the bag is one minus the probability of drawing a black marble. The probability of drawing a black marble is 2 out of 10 So we have 1-2/10=1-1/5=4/5=80%

OpenStudy (zzr0ck3r):

Does this make sense?

OpenStudy (kaywee):

yes thank you !

OpenStudy (zzr0ck3r):

the next one is easier

OpenStudy (kaywee):

im not understanding is the second one 3/10?

OpenStudy (agent0smith):

The probability of drawing a red marble is How many reds? How many total? Divide.

OpenStudy (kaywee):

i got 0.3

OpenStudy (agent0smith):

\[\checkmark \]

OpenStudy (kaywee):

A fair coin is tossed 3 times in a row. What is the probability that heads appears on only the last toss? A. 1/8 B. 3/8 C. 3/16 D. 3/4

OpenStudy (agent0smith):

You might want to list all the possible outcomes. Use H for head and T for tail.

OpenStudy (kaywee):

Someone please help!

OpenStudy (agent0smith):

New questions are better than continually using the same question